# MLLM-FL: Multimodal Large Language Model Assisted Federated Learning on Heterogeneous and Long-tailed Data

### v1

@@ 3.2. Federated Finetuning
$w_{k}^{t+1}\leftarrow w_{k}^{t}-\eta\nabla_{w}L_{\text{loc}}(w^{t};\mathcal{D}^{k})$
@@
(1) $w_{agg}^{t+1}=\sum_{k\in\Omega^{t}}\frac{\left|\mathcal{D}^{k}\right|}{\sum_{k\in\Omega^{t}}\left|\mathcal{D}^{k}\right|}w_{k}^{t+1}$
@@ 3.3. Global Alignment
$w_{agg}^{new}\leftarrow w_{agg}-\eta\nabla_{w}L_{align}(w_{agg};\mathcal{D}_{align})$
@@
$L_{align}=L_{ce}\left(y,p\right)+\beta\cdot KL\left(q\|p\right),$
where $L_{ce}(\cdot,\cdot)$ is the cross-entropy loss. $y$ is the label and $p$ is the output logits vector of the FL models. Since the pretrained CLIP model in the pertaining mechanism has zero-shot image classification capability (Radford et al., 2021), $q$ denotes the output logits vector of the CLIP model. $KL$ is the Kullback-Leibler divergence and $\beta$ is a hyperparameter balanced these two losses.
Table 2. Top-1 classification accuracy(%) on CIFAR-10-LT and CIFAR-100-LT datasets with different FL methods, where th results are referred in (Shang et al., 2022; Shi et al., 2023). The best results are marked in bold.
@@
4. Experiment
4.1. Experiment Setup
@@
We compare MLLM4FL with 13 FL methods: FedAvg (McMahan et al., 2017), FedAvgM (Hsu et al., 2019), FedProx (Li et al., 2018), FedDF (Sui et al., 2020), FedBE (Chen and Chao, 2020), CCVR (Luo et al., 2021) and FedNova (Wang et al., 2020a), Fed-Focal Loss (Sarkar et al., 2020), Ratio Loss (Wang et al., 2020b) and FedAvg with $\tau$-norm (Kang et al., 2019), CReFF (Shang et al., 2022) and CLIP2FL (Shi et al., 2023). The first seven approaches are heterogeneity-oriented , and Fed-Focal Loss, Ratio Loss and FedAvg with $\tau$-norm are imbalance-oriented.
4.2. Experimental Results
Results for CIFAR-10/100-LT are presented in Table 2, where we evaluate the performance of our CLIP2FL against a range of FL approaches on both CIFAR-10-LT and CIFAR-100-LT datasets. Notably, MLLM-FL outperforms other methods in terms of classification accuracy on both datasets. Specifically, at an Imbalance Factor (IF) of 100, which presents a severe imbalance, MLLM shows an improvement of 2.12% and 1.94% in classification accuracy over CLIP2FL for CIFAR-10-LT and CIFAR-100-LT, respectively. Under the condition of IF = 50 or 10, MLLM still manages to enhance performance by around 1%. This underscores MLLM-FL’s effectiveness and its outperforms over competing methods to deal with heterogeous and long-tailed distributions.
Table 3. Top-1 accuracy(%) on ImageNet-LT dataset with different FL method
@@ 4.2. Experimental Results
In the context of ImageNet-LT, Table 3 presents a comparison of the accuracy achieved by our MLLM-FL framework against various FL approaches. The evaluation is segmented into four groups based on the number of samples per class: “Many” (over 100 samples), “Medium” (20 to 100 samples), “Few” (less than 20 samples), and “All” (overall accuracy). While our method may not fully match the performance of CReFF in the “Many” categories, it excels in “Overall” accuracy and the “Medium” and ”Few” category. These results underscore MLLM-FL’s capability not just in enhancing overall model performance but also in significantly improving classification outcomes for categories with fewer samples. The ImageNet-LT results highlight the effectiveness of MLLM-FL in tackling the inherent challenges of long-tailed data distributions.
4.3. Further Analysis
@@
Table 4. Comparison between pretrained and non-pretrained models under constrained training dataset settings. The format is (number of epochs, highest accuracy)
@@
4.3.1. Ablation studies on our pretraining mechanism
@@
To evaluate the effectiveness of pretraining, we conducted a comparative analysis between a pretrained model and a non-pretrained model under a constrained training dataset scenario akin to few-shot learning, aiming to mirror real-world conditions where the amount of data available on each device is limited. We generated subsets of the CIFAR-10 and CIFAR-100 datasets at varying proportions and trained both models for 30 epochs. Our findings, detailed in Table 4, outline the number of epochs required by each model to reach predetermined accuracy thresholds with different training sample sizes, along with the highest accuracy achieved by each model within the 30-epoch span.
@@
Specifically, within the CIFAR-10 context, to attain a target accuracy of $25\%$, the pretrained model consistently outperformed the non-pretrained model, requiring fewer epochs across all sample sizes. Similarly, for the CIFAR-100 dataset, in pursuit of a target accuracy of $15\%$, the pretrained model proved more efficient, also necessitating fewer epochs. Figure 3 further illustrates the superiority of our pretraining approach, demonstrating that models pretrained using our methodology surpass those trained from scratch in terms of accuracy.
@@
Figure 3. The comparative analysis of pretrained and non-pretrained models using 1% subsets of CIFAR-10/100 training data.
[…]
4.3.2. Ablation studies on our global alignment mechanism
In Figure 4, we present the confusion matrix for our model equipped with global alignment within the CIFAR-10-LT dataset, characterized by an imbalance factor of 100. We normalize this confusion matrix by the volume of data in each class. Additionally, we illustrate a normalized confusion matrix for a baseline model devoid of global alignment within our Federated Learning (FL) framework. The visual representations indicate that, with alignment in place, data from each class can be accurately classified. In contrast, the absence of alignment yields inferior results, particularly for classes with few data. Compared to classes with abundant data, those with fewer instances often experience misclassification, with minority class data being inaccurately labeled as belonging to majority classes. This starkly underscores the significance of our global alignment strategy in enhancing both the performance and fairness of the FL system.
@@ 4.3.2. Ablation studies on our global alignment mechanism
Figure 4. The comparative analysis of aligned and non-aligned models with normalized confusion matrices.
[…]
4.3.3. Discussion
Privacy
@@ 4.3.3. Discussion
: At the forefront of federated learning challenges is the protection of user privacy. Our strategy sidesteps the conventional requirement for clients to send gradients back to the server, as seen in methods like CReFF (Shang et al., 2022) and CLIP2FL (Shi et al., 2023). This aspect is vital because the transmission of gradients could enable the server to perform reverse engineering attacks (Geiping et al., 2020; Zhu et al., 2019; Haim et al., 2022), potentially endangering client data confidentiality. By eliminating this step, our method diminishes the likelihood of leaking sensitive client information, promoting a more secure and privacy-centric learning environment.
Computational Efficiency
@@
: Our approach also stands out for its computational economy. Contrary to approaches like CLIP2FL, which necessitate deploying sizable multimodal models such as CLIP on client devices—demanding significant memory and potentially being unfeasible for edge devices with limited resources—our method positions the MLLM solely on the server side. At the client level, we deploy only the compact FL models. This resolution not only addresses memory constraints but also reduces the time and energy expenditure associated with federated local training. Consequently, our framework is rendered more practical and appealing for an extensive array of devices, particularly those with restricted storage capacities.
Compatibility
@@
: Our approach stands out for its adaptability, unlike specific methodologies like CReFF and CLIP2FL that impose unique requirements on federated local training and global aggregation. Our framework can be compatible with a wide array of existing FL algorithms. This includes, but is not limited to, client selection strategies and various techniques aimed at further enhancing client privacy protection. This flexibility ensures that our method can be seamlessly integrated into diverse FL environments and fully leverage the accumulated advancements from previous federated learning research. For a more intuitive comparison, please refer to the Table 1 highlighting the advantages of our method.
5. Conclusion
@@
To overcome the challenges of federated learning in the context of heterogeneous and long-tailed data distributions, we introduced a novel framework, MLLM-FL. This framework is structured around three core stages: global pretraining, federated fine-tuning, and global alignment. This marks the inaugural integration of Multimodal Large Language Models (MLLMs) into an FL system. Leveraging the strong multimodal capacities of MLLMs, our approach taps into the vast yet previously underutilized reservoir of open-source data available online, alongside substantial server-side computational resources. Crucially, our methodology does not compromise privacy nor impose additional computational demands on client devices. Experimental evidence verifies the efficacy of our framework, paving the way for future research to explore a broader array of multimodal tasks beyond image-text interactions, thereby enhancing FL performance across diverse multimodality challenges.

### v2

@@ 3.2 Federated Finetuning
$w_{k}^{t+1}\leftarrow w_{k}^{t}-\eta\nabla_{w}L_{\text{loc}}(w^{t};\mathcal{D}^{k})$ (3)
@@
$w_{agg}^{t+1}=\sum_{k\in\Omega^{t}}\frac{\left|\mathcal{D}^{k}\right|}{\sum_{k\in\Omega^{t}}\left|\mathcal{D}^{k}\right|}w_{k}^{t+1}$ (4)
@@ 3.3 Global Alignment
$w_{agg}^{new}\leftarrow w_{agg}-\eta\nabla_{w}L_{align}(w_{agg};\mathcal{D}_{align})$ (5)
@@
$L_{align}=L_{ce}\left(y,p\right)+\beta\cdot KL\left(q\|p\right),$ (6)
where $L_{ce}(\cdot,\cdot)$ is the cross-entropy loss. $y$ is the label and $p$ is the output logits vector of the FL models. Since the pretrained CLIP model in the pertaining mechanism has zero-shot image classification capability [41], $q$ denotes the output logits vector of the CLIP model. $KL$ is the Kullback-Leibler divergence and $\beta$ is a hyperparameter balanced these two losses.
@@
Table 2: Top-1 classification accuracy(%) on CIFAR-10-LT and CIFAR-100-LT datasets with different FL methods, where th results are referred in [44, 45]. The best results are marked in bold.
@@
4 Experiment
4.1 Experiment Setup
@@
We compare MLLM4FL with 13 FL methods: FedAvg [36], FedAvgM [16], FedProx [29], FedDF [46], FedBE [2], CCVR [34] and FedNova [50], Fed-Focal Loss [43], Ratio Loss [51] and FedAvg with $\tau$-norm [23], CReFF [44] and CLIP2FL [45]. The first seven approaches are heterogeneity-oriented , and Fed-Focal Loss, Ratio Loss and FedAvg with $\tau$-norm are imbalance-oriented.
4.2 Experimental Results
Results for CIFAR-10/100-LT are presented in Tab. 2, where we evaluate the performance of our CLIP2FL against a range of FL approaches on both CIFAR-10-LT and CIFAR-100-LT datasets. Notably, MLLM-LLaVA-FL outperforms other methods in terms of classification accuracy on both datasets. Specifically, at an Imbalance Factor (IF) of 100, which presents a severe imbalance, MLLM shows an improvement of 2.12% and 1.94% in classification accuracy over CLIP2FL for CIFAR-10-LT and CIFAR-100-LT, respectively. Under the condition of IF = 50 or 10, MLLM still manages to enhance performance by around 1%. This underscores MLLM-LLaVA-FL’s effectiveness and its outperforms over competing methods to deal with heterogenous and long-tailed distributions.
@@ 4.2 Experimental Results
Table 3: Top-1 accuracy(%) on ImageNet-LT dataset with different FL method
In the context of ImageNet-LT, Tab. 3 presents a comparison of the accuracy achieved by our MLLM-LLaVA-FL framework against various FL approaches. The evaluation is segmented into four groups based on the number of samples per class: “Many” (over 100 samples), “Medium” (20 to 100 samples), “Few” (less than 20 samples), and “All” (overall accuracy). While our method may not fully match the performance of CReFF in the “Many” categories, it excels in “Overall” accuracy and the “Medium” and "Few" category. These results underscore MLLM-LLaVA-FL’s capability not just in enhancing overall model performance but also in significantly improving classification outcomes for categories with fewer samples. The ImageNet-LT results highlight the effectiveness of MLLM-LLaVA-FL in tackling the inherent challenges of long-tailed data distributions.
4.3 Further Analysis
@@
Table 4: Comparison between pretrained and non-pretrained models under constrained training dataset settings. The format is (number of epochs, highest accuracy)
4.3.1 Ablation studies on our pretraining mechanism
@@
To evaluate the effectiveness of pretraining, we conducted a comparative analysis between a pretrained model and a non-pretrained model under a constrained training dataset scenario akin to few-shot learning, aiming to mirror real-world conditions where the amount of data available on each device is limited. We generated subsets of the CIFAR-10 and CIFAR-100 datasets at varying proportions and trained both models for 30 epochs. Our findings, detailed in Tab. 4, outline the number of epochs required by each model to reach predetermined accuracy thresholds with different training sample sizes, along with the highest accuracy achieved by each model within the 30-epoch span.
@@
Specifically, within the CIFAR-10 context, to attain a target accuracy of $25\%$, the pretrained model consistently outperformed the non-pretrained model, requiring fewer epochs across all sample sizes. Similarly, for the CIFAR-100 dataset, in pursuit of a target accuracy of $15\%$, the pretrained model proved more efficient, also necessitating fewer epochs. Fig. 3 further illustrates the superiority of our pretraining approach, demonstrating that models pretrained using our methodology surpass those trained from scratch in terms of accuracy.
@@
Figure 3: The comparative analysis of pretrained and non-pretrained models using 1% subsets of CIFAR-10/100 training data.
@@
Figure 4: The comparative analysis of aligned and non-aligned models with normalized confusion matrices.
4.3.2 Ablation studies on our global alignment mechanism
In Fig. 4, we present the confusion matrix for our model equipped with global alignment within the CIFAR-10-LT dataset, characterized by an imbalance factor of 100. We normalize this confusion matrix by the volume of data in each class. Additionally, we illustrate a normalized confusion matrix for a baseline model devoid of global alignment within our Federated Learning (FL) framework. The visual representations indicate that, with alignment in place, data from each class can be accurately classified. In contrast, the absence of alignment yields inferior results, particularly for classes with few data. Compared to classes with abundant data, those with fewer instances often experience misclassification, with minority class data being inaccurately labeled as belonging to majority classes. This starkly underscores the significance of our global alignment strategy in enhancing both the performance and fairness of the FL system.
4.3.3 Discussion
Privacy:
@@ 4.3.3 Discussion
Our strategy sidesteps the conventional requirement for clients to send gradients back to the server, as seen in methods like CReFF [44] and CLIP2FL [45]. This aspect is vital because the transmission of gradients could enable the server to perform reverse engineering attacks [12, 72, 13, 47], potentially endangering client data confidentiality. By eliminating this step, our method diminishes the likelihood of leaking sensitive client information, promoting a more secure and privacy-centric learning environment.
Computational Efficiency:
@@
Our approach also stands out for its computational economy. Contrary to approaches like CLIP2FL, which necessitate deploying sizable multimodal models such as CLIP on client devices—demanding significant memory and potentially being unfeasible for edge devices with limited resources—our method positions the MLLM solely on the server side. At the client level, we deploy only the compact FL models. This resolution not only addresses memory constraints but also reduces the time and energy expenditure associated with federated local training. Consequently, our framework is rendered more practical and appealing for an extensive array of devices, particularly those with restricted storage capacities.
Compatibility:
@@
Our approach stands out for its adaptability, unlike specific methodologies like CReFF and CLIP2FL that impose unique requirements on federated local training and global aggregation. Our framework can be compatible with a wide array of existing FL algorithms. This includes but is not limited to, client selection strategies and various techniques aimed at further enhancing client privacy protection. For a more comprehensive comparison, please refer to the Tab. 1 highlighting the advantages of our method.
5 Conclusion
@@
To overcome the challenges of federated learning in the context of heterogeneous and long-tailed data distributions, we introduced a novel framework, MLLM-LLaVA-FL. This framework is structured around three core stages: global pretraining, federated fine-tuning, and global alignment. This marks the inaugural integration of MLLMs into an FL system. Leveraging the strong multimodal capacities of MLLMs, our approach taps into the vast yet previously underutilized reservoir of open-source data available online, alongside substantial server-side computational resources. Crucially, our methodology does not compromise privacy nor impose additional computational demands on client devices. Experimental evidence verifies the efficacy of our framework, paving the way for future research on more multimodal tasks.
Acknowledgement
The research performed at Duke University was supported in part by NSF 2112562 and ARO W911NF-23-2-0224.